For a three-way bulb (60 W, 100 W, 160 W) that is designed to operate at 120 V, find the resistance of each of the two filaments. Assume that the wattage ratings are not limited by significant figures and ignore any heating effects on the resistances.

60 W filament________Ohms
100 W filament__________Ohms
1
Expert's answer
2012-02-14T11:17:42-0500
Power
P=V2/R whence R=V2/P
60W filament 240 Ohms
100W filament 144 Ohms